Instagram surveys 1000 people within the ages 18-44, and this age group makes up 38% of the overall population.

Now, Instagram wants to survey people ages 45-64, who makes up 25% of the overall population. With everything else being equal, how many people should they survey in the 45-64 year age group to get a similar level of accuracy in their estimations for BOTH age groups? Why?
Answer #1

In 18-44 age groups proportion observed, p =0.38

So sample variance is given by p(1-p)/n= 0.38*0.62/1000

Again in the second sample we have p=0.25 so variance will be 0.25*0.75/n where n is the number of people surveyed.

So our required n for same accuracy is obtained by solving the equation

[0.38*0.62/1000]=[0.25*0.75/n]

So n= 1257 (answer)